1.Eight workers can paint an entire house in ten hours. How long would it take five workers to paint the same house?
2. It takes five workers twelve hours to unload one truck. How long would it take six workers to unload the same truck.
3. If twelve workers cna dig a tunnel in 100 days, how long will it take twenty workers to dig the tunnel.

Answer by stanbon(75887)

Amt of time and # of workers are indirectly related:
t = k/n
----
1.Eight workers can paint an entire house in ten hours. How long would it take five workers to paint the same house?
Solve for "k" using "8 workers can paint the house in ten hours"
10 = k/8
k = 80
Equation:
t = 80/n
How log would it take 5 workers to point the house?
t = 80/5 = 16 hrs
============================================
2. It takes five workers twelve hours to unload one truck. How long would it take six workers to unload the same truck.
Find "k":
t = k/n
12 = k/5
k = 60
Equation:
t = 60/n
How long would it take six workers?
t = 60/6 = 10 hrs
========================
3. If twelve workers can dig a tunnel in 100 days, how long will it take twenty workers to dig the tunnel.
100 = k/12
k = 1200
Equation:
t = 1200/n
How long would it take 20 workers?
t = 1200/20 = 60 hrs
=========================
Cheers,
Stan H.
==================